# HYDRA: Model Factorization Framework for Black-Box LLM Personalization

### v2

@@ 1 Introduction
Pre-trained large language models (LLMs) [34, 35, 2, 30, 4, 29] have revolutionized various natural language processing (NLP) tasks, ranging from traditional recommendation systems [61, 12, 8, 6] to modern virtual assistants [9, 23, 62, 65, 43, 64, 46]. Despite their strong capabilities, LLMs require further customization to consistently demonstrate desirable behaviors to each user and achieve optimal performance in specific use cases [47, 42, 60, 10, 13, 14, 24, 56, 57, 44, 58]. As a result, LLM personalization has emerged as a rapidly evolving area of research [20], with the goal of tailoring the emergent abilities of LLMs to meet the unique needs of individual users.
@@ 2 Related Works
In-Context Learning. Vanilla personalized prompting approaches leverage the powerful in-context learning capability of LLMs by using the user’s randomly sampled behavior history as contextual samples. Existing studies [6, 18, 25, 51, 63] have utilized encoding user histories—whether personal ratings, interaction histories, or exemplary reviews—as few-shot examples to facilitate LLMs in generating personalized content in various down-stream applications. Additionally, research has shown that utilizing a longer user history can potentially lead to better performance [5, 63].
@@
Limitations. Similar to in-context learning, personalization based on PAG is prone to being easily distracted by irrelevant information retrieved, especially when there is a shift in user behavior between the current query and the user’s historical records. Despite the improvement upon PAG by retrieving relevant information, RAG-based methods may still suffer from the quality of retrieved information, where the "most relevant" information may not necessarily be the "most useful" information to answer a new query. Additionally, prompting-based methods not only lack deeper personalization analysis due to their reliance on a single centralized model but also lack access to global knowledge due to the user-specific prompt design.
@@ 4.4 Scale-up Analysis
Number of Users in Training. We study the impact of users in the training data, as depicted in Figure 3 (a)-(c). As the number of users increases from 20 to 100, we observe that HYDRA reaches over 90% of best-performance with only 20% of training data. It showcases the effectiveness and stability of HYDRA, with the model performance further enhancing with additional users in training.
@@ 4.5 Empirical Personalization Analysis
Performance under User Behavior Shift. Following the setup in Tan et al. [49] (see Appendix G.3 for additional details), we evaluate HYDRA under user behavior shift when the query is not relevant to all historical records. As illustrated in Figure 5, our proposed HYDRA continues to outperform the current state-of-the-art baselines in the face of user behavior shift. This serves as evidence of its robustness and generalizability to provide a widely applicable solution for black-box LLM personalization. This improvement can be attributed to the personalized reranker, which reevaluates the candidates based on their utility rather than solely on relevance, together with the personalized adapter further adjusts to align with user preferences.
@@ 5 Conclusions
In this paper, we proposed HYDRA, a model factorization framework for black-box LLM personalization that captures and leverages both user-specific and shared behavior patterns from historical data. Compared to prompt-based methods, HYDRA introduces a novel learning-based paradigm for black-box LLM personalization that not only remains highly effective in personalizing outputs by training a reranker with a black-box LLM adapter, but also eliminates the need for access to model parameters, presenting a promising alternative to existing mainstream techniques. HYDRA establishes a foundation for learning-based black-box LLM personalization, facilitating targeted enhancements in user experience and ultimately contributing to the development of human-centric intelligent systems.
@@ G.3 Additional Experimental Details for User Behavior Shift
Following the setup of existing work [49], we utilize an encoder-only language model called DeBERTa-v3 [11] to encode the user’s historical behaviors and the query into high-dimensional representations. We then compute the cosine similarity between the query and all historical items to evaluate their relevance. We select the top 20 items with the lowest relevance scores as irrelevant user history. As depicted in Figure 5, our proposed HYDRA still outperforms the state-of-the-art baselines. This performance improvement may stem from the personalized reranker, which reranks the candidates based on their usefulness rather than just relevance. The personalized adapter then further adapts the model’s generations to align with user preferences. In addition, prompt-based methods tend to underperform when there is a distribution shift between the few-shot demonstrations and user queries. This issue can become even more severe under personalized scenarios, where there might be a mismatch between the selected user history and the user query.
@@ G.5 Effect of Black-Box LLM Adapters
We evaluate the effectiveness of different black-box LLM adapters in Table 7. Due to the inaccessibility to model internal parameters, BBox-Adapter is the only black-box LLM adapter available, apart from our proposed HYDRA-Adapter. We observe that HYDRA-Adapter outperforms or achieves competitive performance compared to BBox-Adapter when applied under the same personalization settings (as discussed in Section 3.4). This result demonstrates the effectiveness of our adapter module design, which captures both global and user-specific knowledge to further adapt to personalized model outputs.

### v3

@@ 1 Introduction
Pre-trained large language models (LLMs) [36, 37, 2, 32, 5, 31] have revolutionized various natural language processing (NLP) tasks, ranging from traditional recommendation systems [63, 13, 9, 7] to modern virtual assistants [10, 24, 64, 67, 45, 66, 48]. Despite their strong capabilities, LLMs require further customization to consistently demonstrate desirable behaviors to each user and achieve optimal performance in specific use cases [49, 44, 62, 11, 14, 15, 25, 58, 59, 46, 60, 4]. As a result, LLM personalization has emerged as a rapidly evolving area of research [21], with the goal of tailoring the emergent abilities of LLMs to meet the unique needs of individual users.
@@ 2 Related Works
In-Context Learning. Vanilla personalized prompting approaches leverage the powerful in-context learning capability of LLMs by using the user’s randomly sampled behavior history as contextual samples. Existing studies [7, 19, 26, 53, 65] have utilized encoding user histories-whether personal ratings, interaction histories, or exemplary reviews-as few-shot examples to facilitate LLMs in generating personalized content in various down-stream applications. Additionally, research has shown that utilizing a longer user history can potentially lead to better performance [6, 65].
@@
Limitations. Similar to in-context learning, personalization based on PAG is prone to be easily distracted by irrelevant information retrieved, especially when there is a shift in user behavior between the current query and the user’s historical records. Despite the improvement upon PAG by retrieving relevant information, RAG-based methods may still suffer from the quality of retrieved information, where the "most relevant" information may not necessarily be the "most useful" information to answer a new query. Additionally, prompting-based methods not only lack deeper personalization analysis due to their reliance on a single centralized model but also lack access to global knowledge due to the user-specific prompt design.
@@ 4.4 Scale-up Analysis
Number of Users in Training. We study the impact of users in the training data, as depicted in Figure 3 (a)-(c). As the number of users increases from 20 to 100, we observe that HYDRA reaches over 90%
@@
Number of Users in Training Set. We also conduct additional scale-up experiments (Table 3) to evaluate HYDRA with an increased number of users, increasing from 100 to 1000, across all five tasks. Our findings from the scale-up experiments show that HYDRA maintains its performance advantages over baselines as the number of users increases.
Dataset ($\rightarrow$) LaMP-2N LaMP-2M LaMP-3 LaMP-4 LaMP-5
gpt-3.5-turbo 0.638 0.499 0.412 0.347 0.540 0.851 0.133 0.119 1.043 0.439 0.371 6.018
ICL-Random (k=1) 0.598 0.476 0.392 0.335 0.676 0.959 0.147 0.132 1.330 0.457 0.396 8.118
ICL-Random (k=2) 0.632 0.499 0.376 0.311 0.562 0.871 0.151 0.137 2.388 0.451 0.393 8.550
ICL-Random (k=4) 0.630 0.518 0.392 0.352 0.440 0.740 0.161 0.146 2.418 0.457 0.396 8.404
RAG (k=1) 0.610 0.486 0.408 0.345 0.602 0.871 0.154 0.138 1.649 0.468 0.405 7.820
RAG (k=2) 0.624 0.479 0.380 0.315 0.559 0.836 0.161 0.149 2.958 0.480 0.419 9.021
RAG (k=4) 0.656 0.524 0.392 0.339 0.391 0.716 0.167 0.155 3.615 0.479 0.418 9.108
PAG (k=0) 0.618 0.489 0.404 0.340 0.583 0.872 0.161 0.141 1.950 0.460 0.405 7.372
PAG (k=1) 0.630 0.500 0.418 0.357 0.414 0.787 0.163 0.153 2.934 0.474 0.414 8.372
HYDRA 0.748 0.551 0.446 0.373 0.328 0.656 0.175 0.167 4.772 0.508 0.442 9.519
Table 3: Scale-up experiment results on the LaMP benchmark, including $1000$ and $500$ users during training and testing, respectively.
@@ 4.5 Empirical Personalization Analysis
Performance under User Behavior Shift. Following the setup in Tan et al. [51] (see Appendix G.3 for additional details), we evaluate HYDRA under user behavior shift when the query is not relevant to all historical records. As illustrated in Figure 5, our proposed HYDRA continues to outperform the current state-of-the-art baselines in the face of user behavior shift. This serves as evidence of its robustness and generalizability in providing a widely applicable solution for black-box LLM personalization. This improvement can be attributed to the personalized reranker, which reevaluates the candidates based on their utility rather than solely on relevance, together with the personalized adapter further adjusts to align with user preferences.
Figure 6: Experiments on LaMP benchmark of significant disparities in user behavior history.
Effect of Significant Disparities in User Behavior. To consider more extreme cases of disparities in user behavior, we retrain HYDRA on a mixture of 50% users with the fewest interactions and another 50% users with the most interactions (blue in Figure 6, compared to the random selection in orange). The black and gray dashed lines represent the best-performing baselines on the first and second metrics. The experimental results demonstrate that HYDRA consistently outperforms existing baselines, even under extreme cases. Compared to the previous random selection of training users, HYDRA achieves relatively lower performance due to the imbalance of training samples for dense users and sparse users. By leveraging the global information in shared parameters, knowledge can be effectively transferred from dense users to sparse users, thereby enabling further personalization through the utilization of sparse user-specific head models.
@@ 5 Conclusions
In this paper, we proposed HYDRA, a model factorization framework for black-box LLM personalization that captures and leverages both user-specific and shared behavior patterns from historical data. Compared to prompt-based methods, HYDRA introduces a novel learning-based paradigm for black-box LLM personalization that not only remains highly effective in personalizing outputs by training a reranker with a black-box LLM adapter, but also eliminates the need for access to model parameters, presenting a promising alternative to existing mainstream techniques. Experimental results demonstrate that HYDRA outperforms state-of-the-art RAG and PAG baselines by an average relative improvement of 9.01% across five diverse personalization tasks. HYDRA establishes a foundation for learning-based black-box LLM personalization, facilitating targeted enhancements in user experience and ultimately contributing to the development of human-centric intelligent systems.
Acknowledgments and Disclosure of Funding
This work was supported by NSF grants IIS-2008334, IIS-2106961, CAREER IIS-2144338, IIS-2403240, and Azure credits from the Microsoft Accelerating Foundation Models Research Program.
@@ G.3 Additional Experimental Details for User Behavior Shift
Following the setup of existing work [51], we utilize an encoder-only language model called DeBERTa-v3 [12] to encode the user’s historical behaviors and the query into high-dimensional representations. We then compute the cosine similarity between the query and all historical items to evaluate their relevance. We select the top 20 items with the lowest relevance scores as irrelevant user history. As depicted in Figure 5, our proposed HYDRA still outperforms the state-of-the-art baselines. This performance improvement may stem from the personalized reranker, which reranks the candidates based on their usefulness rather than just relevance. Additionally, the personalized adapter also adapts the model’s generations to align with user preferences. In addition, prompt-based methods tend to underperform when there is a distribution shift between the few-shot demonstrations and user queries. This issue can become even more severe under personalized scenarios, where there might be a mismatch between the selected user history and the user query.
@@ G.5 Effect of Black-Box LLM Adapters
We evaluate the effectiveness of different black-box LLM adapters in Table 8. Due to the inaccessibility to model internal parameters, BBox-Adapter is the only black-box LLM adapter available, apart from our proposed HYDRA-Adapter. Based on Table 8, it can be observed that HYDRA-Adapter outperforms BBox-Adapter when applied under the same personalization settings (as discussed in Section 3.4). This result demonstrates the effectiveness of our adapter module design, which captures both global and user-specific knowledge to further adapt to personalized model outputs.
@@
G.6 Additional Datasets
Dataset ($\rightarrow$) MovieLens-1M Recipe
Method ($\downarrow$) MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$
gpt-3.5-turbo 0.780 1.208 0.880 1.149
ICL-Random (k=1) 0.880 1.2649 0.960 1.233
ICL-Random (k=2) 0.800 1.095 0.980 1.192
ICL-Random (k=4) 0.820 1.225 0.760 1.114
RAG (k=1) 0.800 1.166 0.980 1.192
RAG (k=2) 0.700 1.030 0.880 1.114
RAG (k=4) 0.640 1.000 0.820 1.049
PAG (k=0) 0.820 1.147 0.860 1.128
PAG (k=1) 0.760 1.010 0.800 1.063
HYDRA 0.600 0.908 0.720 0.938
Table 9: Experiments on two additional personalization datasets, MovieLens-1M and Recipe, focusing on predicting users’ personal ratings for movies or recipes based on their historical rating patterns.
To further validate the effectiveness of HYDRA, we conduct extensive experiments on two additional widely used personalization datasets [29], MovieLens-1M and Recipe, focusing on predicting users’ personal ratings for movies or recipes based on their historical rating patterns. The experimental results (Table 9) demonstrate that HYDRA outperforms the best-performing baselines by 8.0% on Movielens and 10.3% on Recipe, respectively.
G.7 Efficiency Analysis
Assume that we have $N_{\text{train}}$ users in the training data and $N_{\text{test}}$ users in the test set. We adopt the transformer architecture, specifically the Longformer-base, as the reranker and the adapter base models. Consequently, the time complexity of all stages should be proportional to that of the attention mechanism in transformers, $O(dL^{2})$, where $L$ indicates the sequence length and $d$ indicates the hidden dimension. Additionally, the training process will go through the transformer for $T$ epochs, while inference only requires one. For each user in the HYDRA-Reranker training data, we augment $M$ random historical records from the user’s corresponding profile. The retriever then retrieves the top-$M$ ($M=20$ by default) relevant historical records to form training samples. Thus, for each user, we collect $M^{2}+1$ training samples. For each user in the HYPER-Adapter training data, we consider all the user historical records. For each record, we leverage model randomness to generate $k$ ($k=8$ by default) samples for the adapter to select. Consequently, we can have $\bar{H}k$ training samples per user, where $\bar{H}$ denotes the average number of histories per user.
Method Mode Time Complexity LaMP-2N LaMP-2M LaMP-3 LaMP-4 LaMP-5
HYDRA-Reranker Training $O(N_{\text{train}}(M^{2}+1)TL^{2}d)$ 31m10s 41m51s 50m37s 1h1m31s 1h8m16s
HYDRA-Reranker Fit New User $O(N_{\text{test}}(M^{2}+1)TL^{2}d)$ 18m8s 21m17s 25m36s 33m52s 31m25s
HYDRA-Reranker Inference $O(N_{\text{test}}L^{2}d)$ 3m4s 3m1s 3m7s 4m38s 5m14s
HYDRA-Adapter Training $O(N_{\text{train}}k\bar{H}TL^{2}d)$ 1h10m17s 2h2m16s 2h1m59s 3h56m47s 3h19m42s
HYDRA-Adapter Fit New User $O(N_{\text{test}}k\bar{H}TL^{2}d)$ 28m15s 1h7m27s 1h19s 2h23m10s 1h59m2s
HYDRA-Adapter Inference $O(N_{\text{test}}kL^{2}d)$ 4m16s 4m17s 4m17s 5m53s 5m59s
Table 10: Time complexity analysis with running time summary on the LaMP benchmark.
@@ Appendix H Case Studies
H.1 Case Studies for Effectiveness of Reranker
Table 11 illustrates two observed cases during the reranking process. In Case 1, the retriever accurately retrieves the text in the optimal order, meaning that the top outputs are the most relevant to the target. In this scenario, the reranker successfully generates scores that align with the retriever, preserving the correct relevance. In Case 2, when the retriever fails to retrieve the text in the optimal order by relevance, the reranker assigns higher scores to the more relevant terms, thereby mitigating the incorrect retrieval ordering. This emphasizes the significance of HYDRA-reranker by offering an additional measurement dimension that is dedicated to assessing usefulness.